(3) Solve the IVP + 6y(t) + 9 Sy()dt = 1, y(0) = 0. (4) Find a(t) that satisfies e(t) = e-t +S* sinh(t – 7)2(7) dt.
Answer #1

+66(0)4 96 y (edt = 1 dt t=0 © Solution 4 + 6y (4) #9 6 46/4 = 16 - yo)=0 from 0 dy - y(o)= 1 Differentiating equation with&(t) = et + Steinh (t-1) dt I thash (+-) alt) - Losh (t-o) (0) -cosh (A) (+) = ēt +S4G) - ēt] +(+) -105 () &(t) = axit) -